A pilot flying at 100km/h along horizontal level 300m above the ground wishes to bomb a tank travelling at 30km/h in the same direction on the ground. How many seconds before he is directly above the tank must he release the bomb and what must be the distance measured in the horizontal direction separating the plane and the tank?

Grade:11

1 Answers

Venkat
5 years ago
A pilot flying at 100km/h along horizontal level 300m above the ground wishes to bomb a tank travelling at 30km/h in the.....
 
Pilot’s net velocity = 100-30 = 70 km/h x 1000/3600  = 19.44 m/s
For a projectile